a body falls freely from rest and has velocity v after it falls through height h the distance it has to fall down father from its velocity to become double

Grade:10

1 Answers

Arun
6 years ago
For first 2 m,
v^2 = 0^2 + 2g*h
v^2 = 2gh
After that let it traverse d m futher more when its v becomes double .i.e. 2v
(2v)^2 = v^2 + 2gd
2gd = 3v^2
d = 3v^2/2g
put the value of v^2
d = 3*2gh/2g = 3h meter.
 
Regards
Arun (askIITians forum expert)